OpenStudy (anonymous):

Multiply: (x + 1)(x - 2).

myininaya (myininaya):

\[x(x-2)+1(x-2)\]

myininaya (myininaya):

I use the distributive property twice

myininaya (myininaya):

now you do it twice again

OpenStudy (anonymous):

Each value in each group should be multiplied by each value in the other group: \[(x \times x) + (x \times -2) + (1 \times x) + (1 \times -1)\]

OpenStudy (anonymous):

I think it's x^2 - 2x - 2

myininaya (myininaya):

\[-2x+1x=(-2+1)x \neq -2x\]

OpenStudy (anonymous):

What?

myininaya (myininaya):

(-2+1) is not -2

OpenStudy (anonymous):

x^2-2

myininaya (myininaya):

\[x(x-2)+1(x-2)=(x^2-2x)+(x-2)=x^2+x(-2+1)-2\]

OpenStudy (anonymous):

2x-2

OpenStudy (anonymous):

You can use the FOIL method (x + 1)(x - 2) Use the FOIL method to get the product F = Multiply the First numbers x * x = x^2 O= Multiply the Outer numbers x * -2 = -2x I = Multiply the Inner numbers 1 * x = x L = Multiply the Last numbers 1 * -2 = -2 Now put it together x^2 - 2x +x -2 Combine like terms x^2 - x -2

OpenStudy (anonymous):

Thanks! :)